Which of the following six functions represent a travelling wave pulse whose shape does not change with time?

  1. 2e-5(x-300t)2
  2. (2t-5x)3
  3. 9x2-42xt+49t2
  4. x2-49t2
  5. sqrt[(3t-5x)2+5]
  6. (5+2x-3t)/(5+2x+3t)
1
Expert's answer
2022-01-12T08:31:20-0500

Solution

For checking six functions represent a travelling wave pulse whose shape does not change with time

"\frac{ ∂^2\psi}{ ∂x^2}=\frac{1}{v^2}\frac{ ∂^2\psi}{ ∂t^2}" ............ (1)


Now function

"\psi=2e^{{-5}}{(x-300t) ^2}"

"\psi=2e^{{-5}}{(x-300tx+t^2)}"

Then it's partial derivatives

"\frac{ ∂^2\psi}{ ∂x^2}=4e^{-5}"

And also

"\frac{ ∂^2\psi}{ ∂t^2}=4e^{-5}"

Then putting in equation (1)

Then velocity is found as fullfilled condition.


Now other

"\psi=9x^2-42xt+49t^2"

Then partial derivatives

"\frac{ ∂^2\psi}{ ∂x^2}=18"

And

"\frac{ ∂^2\psi}{ ∂t^2}=98"

Whicha also satisfy equation.


So this is answer.